OpenStudy (anonymous):

Sin780sin480 +cos120cos60=0.5 Prove the above without using the trigonometry table.

OpenStudy (campbell_st):

well sin 780 = sin(60) sin480 = sin(60) cos(120) = - cos(60) using exact values it  \[\frac{\sqrt{3}}{2}\times \frac{\sqrt{3}}{2} + \frac{-1}{2} \times \frac{1}{2}\] now just simplify

OpenStudy (anonymous):

Ya bug the problem specifies that solve without using trigonometric values.......

OpenStudy (campbell_st):

it says without using tables... the values shown are exact values... and are known by most students who study trig.

OpenStudy (anonymous):

I assumed that the table meant the normal trigonometric table....... That shows the exact value.... Got it cleared.... Thank you..... :) :)